OpenStudy (anonymous):

A family spends 1/10 of its annual income for housing, 1/4 for food and clothing, 1/5 for general expenses, and 2/15 for entertainment. What fractional part of their income is spent on these items altogether?

OpenStudy (anonymous):

You're adding together \[\frac{1}{10}+\frac{1}{4}+\frac{1}{5}+\frac{2}{15}\]Find the lowest common multiple of the denominators first of all.

OpenStudy (anonymous):

k